The average American gets a haircut every 42 days. Is the average smaller for college students? The data below shows the results of a survey of 11 college students asking them how many days elapse between haircuts. Assume that the distribution of the population is normal.

 44, 31, 38, 48, 44, 38, 38, 44, 38, 44, 38 


What can be concluded at the α = 0.10 level of significance? 

a. For this study, we should use

b. The null and alternative hypotheses would be: 

c. The test statistic t

d. The p-value = 

e. The p-value is 

f. Based on this, we should _______ the null hypothesis
Answer #1

Ans:

sample mean=40.455

n=11

sample standard deviation=4.762

a)t- test for a population mean

b)

Ho : μ = 42

Ha : μ < 42

c)

Test statistic:

t=(40.455-42)/(4.762/SQRT(11))

t=-1.076

df=11-1=10

d)p-value=tdist(1.076,10,1)=0.154

e)p-value is greater than alpha.

f)Fail to reject the null hypothesis.